# Abdoul Kader Kaboré

### v3

@@ 4 Experimental Results
These results contrast with prior work (Siddiq et al., 2024a), which reports higher compilation rates due to method-level test generation. Class-level dependencies and inter-method relationships add complexity, increasing failures as LLMs struggle to resolve imports, hierarchies, and scopes.
Finding 7: The gap between syntactical correctness and compilability remains substantial—only 7.2% of all test suites and 12% of syntactically correct ones successfully compile. Class-level dependencies introduce additional complexity, making compilation significantly harder than the method-level approach used in prior studies.
@@
Finding 8: LLM hallucinations are a major source of compilation errors. CMD is dominated by hallucinated missing symbols, Defects4J struggles with encapsulation violations, and SF110 suffers from redundant test generation.
Prompt Engineering and Hallucination Mitigation. Figure LABEL:fig:compilation-error-heatmap reveals the distribution of non-zero error types across different machine learning models and prompt engineering techniques, with the y-axis dynamically adapting to display error types specific to each model and prompt configuration. The heatmap (Figure LABEL:fig:compilation-error-heatmap) shows CoT and GToT reduce hallucinations more effectively than ZSL and FSL. CoT lowers ’Private Access Issue’ (PAI) errors to 2.96% (GPT-3.5) and 1.50% (GPT-4) in CMD, improving encapsulation adherence. GToT nearly eliminates errors in Defects4J through structured dependencies. However, hallucinations in symbol resolution and redundant test generation persist despite structured reasoning.
@@
Finding 9: Reasoning-based prompting (CoT, GToT) significantly reduces errors, particularly Private Access Issues, whereas Zero-shot Learning exacerbates compilation failures by generating unstructured, non-compilable test cases.
4.3 [RQ3]: Static Analysis for Quality.
[Experiment Goal]: This research question evaluates LLM-generated test suite quality through static analysis of coding standard compliance, complexity, and potential defects. We assess adherence to Java conventions, maintainability based on complexity metrics, and structural flaws to determine practical usability.
[Experiment Design]: We evaluate test quality using three approaches. Checkstyle [^7] https://checkstyle.sourceforge.io assesses adherence to Google Java StyleCheckstyle [^8] https://google.github.io/styleguide/javaguide.html and Sun Code Conventions [^9] https://www.oracle.com/java/technologies/javase/codeconventions-contents.html, measuring formatting consistency and standard violations. PMD [^10] https://pmd.github.io/ analyzes cyclomatic complexity (independent paths) and cognitive complexity (comprehension ease) (Dantas and Maia, 2021). SpotBugs [^11] https://pmd.github.io/ detects potential defects categorized by severity: Scariest (critical issues), Scary (high-risk errors), Troubling (moderate concerns), and Of Concern (minor warnings). We compared prompting techniques (ZSL, FSL, CoT, ToT, GToT) across SF110, Defects4J, and CMD datasets.
@@
Coding Standard Compliance. LLM-generated test suites show varying coding standard adherence (Figure LABEL:fig:style-heatmap). FSL achieves highest compliance with Google Java Style and Sun Code Conventions, producing fewer violations overall.
For Google Java Style (Figure LABEL:fig:google-heatmap), indentation violations dominate across all prompting techniques, with GToT showing the lowest count (1757.5 vs. 2305.9 for ZSL). However, GToT increases import-related violations (140.6 CustomImportOrder and 93.1 AvoidStarImport violations) compared to FSL (71.3 and 23.4 respectively). For Sun Code Conventions (Figure LABEL:fig:sun-heatmap), MissingJavadocMethod violations are highest with CoT (566.5) and GToT (564.0), while GToT significantly reduces MagicNumber violations (134.8 vs. 323.4 for ZSL). These patterns suggest that reasoning-based prompts (CoT, ToT, GToT) improve code structure but sacrifice documentation and import organization. FSL strikes the best balance, minimizing both structural and documentation violations. These findings partially align with  Tang et al. (2024), who reported high ZSL violations, though our results show FSL and CoT significantly improve compliance while ToT and GToT produce tests with better logical structure but less style adherence.
Finding 10: FSL achieves the highest coding standard compliance, followed by CoT, which also reduces violations and improves adherence to style guidelines. In contrast, GToT enhances logical organization but introduces more violations, reinforcing the trade-off between structured prompting and strict adherence to coding conventions.
Code Complexity Analysis. Table 9 shows most LLM-generated test suites exhibit low complexity across prompting strategies. GToT produces the highest proportion (99.99%) of low-cyclomatic-complexity tests while generating no instances of high cognitive complexity, demonstrating its ability to create structured yet maintainable and understandable test cases.
@@
These findings align with Tang et al. (2024), who found ChatGPT tests under ZSL exhibit low complexity. However, our results show reasoning-based prompting (CoT, ToT, GToT) refines logical structure while maintaining this low complexity. While ZSL produces simple tests, it lacks the structured refinement of guided reasoning approaches, highlighting the advantage of stepwise logical improvements in test generation.
Finding 11: GToT produces simpler, more maintainable test suites with the lowest cyclomatic and cognitive complexity. ZSL and CoT also generate structurally sound tests, while FSL primarily improves coding standard compliance without significantly altering complexity.
Static Analysis with SpotBugs. Table 10 reveals 8,461 issues in 20,530 compilable tests (55.58% minor ”Of Concern” warnings). FSL generated the most Scariest issues (61.72%), while GToT produced the most Scary (63.92%) and Troubling (69.10%) defects. ZSL had fewer severe issues but simpler tests. This highlights a key trade-off: structured reasoning improves organization but increases subtle defects. Unlike Tang et al. (2024)’s findings with ZSL, our study shows FSL and GToT’s enhanced structure comes with more severe defects, requiring better balance between structured reasoning and defect mitigation.
@@
ZSL = Zero-shot Learning, FSL = Few-shot Learning, CoT = Chain-of-Thought, ToT = Tree-of-Thought, GToT = Guided Tree-of-Thought.
Finding 12: FSL introduces the most critical bugs, while ZSL produces the safest but least functional test cases. GToT improves test structure but increases subtle defects, requiring additional validation.
@@
[Experiment Goal]: This experiment examines whether LLM-generated test suites meet developer readability expectations compared to EvoSuite. We focus on clarity, maintainability, and formatting. Using an automated readability model correlated with human evaluation, we assess how prompt engineering impacts readability compared to EvoSuite tests.
[Experiment Design]: We employ Scalabrino et al. (2018) readability model, trained on Java code including JUnit tests. This model evaluates syntactic structure, identifier clarity, textual coherence, and structural readability metrics, providing assessment beyond formatting-based analysis. We evaluate across three datasets (SF110, Defects4J, CMD) using multiple prompting techniques (ZSL, FSL, CoT, ToT, GToT), extracting readability scores for each test suite and aggregating results by model and strategy. Higher scores indicate more structured, easier-to-interpret tests—essential for real-world development. We also compare score distributions to identify approaches yielding improved structure and reduced complexity.
@@
Dataset-Level Readability Analysis. Table 11 shows LLM-generated test suites consistently achieve higher readability than EvoSuite across all datasets. Readability varies with codebase complexity: Defects4J exhibits highest readability scores (GPT-3.5-Turbo: 81.73%, Mistral 7B: 69.06% vs. EvoSuite: 57.33%), followed by SF110 (GPT-3.5-Turbo: 75.95%, Mistral 7B: 70.28% vs. EvoSuite: 50.67%). CMD presents lowest scores (GPT-3.5-Turbo: 60.93%, Mistral 7B: 57.13%), indicating modern, evolving codebases challenge LLMs more. While LLMs consistently outperform EvoSuite in readability, their effectiveness varies with dataset complexity.
Finding 13: LLM-generated tests outperform EvoSuite in readability. However, readability varies with dataset complexity, as shown in Table 11. CMD, which contains modern and evolving codebases, presents the greatest challenge, yielding the lowest readability scores across all models.
@@
LLM vs. EvoSuite: Readability Performance. Table 12 shows LLM-generated tests consistently surpass EvoSuite in readability. For GPT-3.5-Turbo, GToT achieves the highest score (84.75% vs. EvoSuite’s 50.77%), followed by CoT (82.77%) and ToT (73.32%), while ZSL (77.93%) and FSL (73.49%) score lower. Similarly, for Mistral 7B, GToT leads (76.35%), followed by ToT (69.64%), with ZSL (63.79%) and FSL (60.40%) producing lowest scores. These results demonstrate reasoning-based prompting’s importance for readability.
Finding 14: LLM-generated test suites improve readability over EvoSuite by 21–40%, as shown in Table 12. This highlights the significant advantage of LLMs in producing more comprehensible test code.
@@
Prompting Strategies: Influence on Readability. Table 13 shows reasoning-based prompting significantly enhances test suite clarity. GToT and CoT consistently achieve highest readability scores across datasets. For GPT-3.5-Turbo on CMD, CoT leads (67.99%), followed by GToT (61.86%), while ZSL (57.67%) and FSL (55.76%) score lower. Similarly, for GPT-4 on CMD, GToT achieves highest readability (67.71%). These results confirm minimal prompting provides insufficient guidance for well-structured tests, while reasoning-based approaches are essential for producing maintainable test suites.
@@
Finding 15: CoT and GToT consistently enhance readability, even on complex datasets like CMD, where unstructured prompting results in lower scores.
Model-Specific Readability Trends. Table 13 reveals how different LLMs respond to reasoning-based prompting. GPT-3.5-Turbo performs best with CoT (67.99%), while GPT-4 excels with GToT (67.71%), confirming larger models benefit from structured reasoning. Mistral 7B shows higher variability, reaching 64.61% with GToT, indicating sensitivity to prompt design. Mixtral 8x7B performs exceptionally with ToT (79.93%), though its limited sample size affects stability.
Finding 16: Model architecture, parameter size, and prompting strategy significantly impact readability, with larger models consistently improving when guided by reasoning-based prompting.
@@
[Experiment Design]: We use JaCoCo [^12] https://www.jacoco.org/jacoco/ to measure three metrics: line coverage (executed source code lines), instruction coverage (executed bytecode instructions), and method coverage (methods invoked at least once). Following Tang et al. (2024), we exclude branch coverage as it only considers conditional branches and often yields missing values. We evaluate all compilable test cases across five prompting strategies (ZSL, FSL, CoT, ToT, GToT) against EvoSuite-generated tests across SF110, Defects4J, and CMD datasets. We compute statistical measures (min, max, quartiles, median, mean) to comprehensively analyze structural coverage.
@@
EvoSuite vs. LLMs: Overall Coverage Performance. Tables 14, 15, and 16 reveal a significant performance gap between LLMs and EvoSuite. EvoSuite consistently outperforms LLMs across all metrics, achieving median coverage of 94.34% (line), 95.0% (instruction), and 100% (method) through search-based optimization. LLM-generated tests show higher variability, with reasoning-based prompting significantly improving coverage over ZSL and FSL. GPT-3.5-Turbo with CoT achieves the highest LLM median coverage (75.47% line, 76.0% instruction, 100% method), though all LLM configurations fall short of EvoSuite’s systematic exploration capabilities.
@@
Finding 17: EvoSuite outperforms all LLMs in coverage metrics. However, reasoning-based techniques (CoT, GToT, ToT) significantly enhance LLM test effectiveness, reducing the performance gap.
Dataset-Level Coverage Trends. Tables 1415, and 16 show coverage varies significantly across datasets. Defects4J achieves highest coverage (GPT-3.5-Turbo: 82.74% line, 85.0% instruction, 100% method), suggesting its structured design facilitates test generation with clearer functional boundaries. In contrast, SF110 reports much lower coverage (3.57% line, 2.0% instruction, 16.67% method) due to higher complexity and diverse architectures. CMD reports highest median coverage (100% for some prompts), though this likely reflects its smaller size and modular structure rather than superior LLM capabilities. The contrast between datasets highlights structural influences: Defects4J benefits from potential training exposure (Sallou et al. (2023)), clearer functional boundaries, and modularized code, while SF110 presents challenges through higher complexity and lower public accessibility. Finding 18: LLMs achieve higher coverage on well-documented datasets like Defects4J but struggle with complex projects like SF110, where greater structural variability makes test generation more challenging. CMD’s high coverage is likely due to its small, modular dataset rather than an inherent LLM advantage.
@@
Finding 19: Dataset characteristics influence LLM performance—Defects4J’s high coverage likely stems from pretraining exposure while SF110’s limited representation restricts generalization, highlighting the need for careful dataset selection to mitigate data leakage in test generation studies.
Impact of Reasoning-Based Prompting on Coverage. Tables 141516 show reasoning-based techniques (CoT, GToT, ToT) consistently improve coverage compared to ZSL and FSL. For line coverage, GPT-3.5-Turbo with CoT achieves highest median (75.47%), while ZSL and FSL produce significantly lower coverage (27.68% and 16.14%). Mistral 7B performs best with GToT (30.0%), though still below GPT-3.5-Turbo. For instruction coverage, EvoSuite maintains advantage (95.0%). Among LLMs, GPT-3.5-Turbo with CoT (76.0%) and GToT (45.0%) lead, while Mistral 7B’s best result with GToT (26.0%) indicates smaller models’ limitations. For method coverage, GPT-3.5-Turbo with CoT reaches 100% in some cases, while Mistral 7B improves with GToT (89.59%) but still lags. These results demonstrate reasoning-based prompting enhances LLM performance significantly but doesn’t fully bridge the gap with EvoSuite.

### v4

@@ 4 Experimental Results
These results contrast with prior work (Siddiq et al., 2024a), which reports substantially higher compilation rates. A key reason lies in the different experimental setups. Siddiq et al. (2024a) adopt a method-level perspective, generating one test class per method under test, with all necessary imports explicitly included and post-processing heuristics applied to repair non-compilable outputs. In contrast, our study operates at the class level, where each generation targets the full class under test, often containing multiple interdependent methods and relying solely on the code available in the dataset without import injection or post-repair. Furthermore, our benchmarks (Defects4J, CMD, SF110) encompass larger and more heterogeneous projects than HumanEval or the filtered SF110 subset used by Siddiq et al. (2024a), which further contributes to differences in compilation outcomes.
Finding 9: The gap between syntactical correctness and compilability remains substantial, only 7.2% of all generated test suites and 12% of syntactically correct ones successfully compile. Compared to prior work, which reports higher compilability when generating one test class per individual method under test with injected imports and repair heuristics, our study evaluates entire classes under test (including all their interdependent methods) without import injection or post-repair. This setup highlights the additional challenges LLMs face when handling complete classes rather than isolated methods.
LLMs vs. EvoSuite on Syntactic Correctness and Compilability. Tables 8 and 9 allow a direct comparison between LLMs and EvoSuite. EvoSuite exhibits near-perfect reliability, with all generated test suites being syntactically correct and only 210 non-compilable cases out of 11,100 in Defects4J ($\approx$1.9%), while achieving 100% compilability on SF110. By contrast, LLMs display a much higher fragility: across datasets, only 89.81% of code-containing outputs were syntactically correct, and compilability drops drastically, with success rates below 10% in most cases. Reasoning-based prompting (e.g., GToT) mitigates syntax issues and improves compilability, but still falls far short of EvoSuite’s stability. This contrast underscores that while LLMs can approach EvoSuite in raw generation capacity (RQ0), their outputs are substantially less reliable when measured by structural validity and executable quality.
Finding 10: Compared to EvoSuite, LLMs remain substantially weaker in syntactic correctness and compilability: while EvoSuite compiles nearly all its generated tests (with only $\approx$1.9% failures in Defects4J), LLMs succeed on fewer than 10% of their syntactically correct outputs. This highlights that prompt engineering improves LLM reliability but does not yet close the gap with mature SBST tools.
@@
Finding 11: LLM hallucinations are a major source of compilation errors. CMD is dominated by hallucinated missing symbols, Defects4J struggles with encapsulation violations, and SF110 suffers from redundant test generation.
Prompt Engineering and Hallucination Mitigation.
Figure 5 reveals the distribution of non-zero error types across different machine learning models and prompt engineering techniques, with the y-axis dynamically adapting to display error types specific to each model and prompt configuration. The heatmap (Figure 5) shows CoT and GToT reduce hallucinations more effectively than ZSL and FSL. CoT lowers ’Private Access Issue’ (PAI) errors to 2.96% (GPT-3.5) and 1.50% (GPT-4) in CMD, improving encapsulation adherence. GToT nearly eliminates errors in Defects4J through structured dependencies. However, hallucinations in symbol resolution and redundant test generation persist despite structured reasoning.
(a) By GPT 3.5-Turbo and by prompt
(b) By Mistral 7B and by prompt engineering
(c) By GPT 4 and by prompt
(d) By Mixtral 8x7B and by prompt engineering
Figure 5: Heatmap of compilation errors by model and prompt engineering.
@@
Finding 12: Reasoning-based prompting (CoT, GToT) significantly reduces errors, particularly Private Access Issues, whereas Zero-shot Learning exacerbates compilation failures by generating unstructured, non-compilable test cases.
Relation to prior work.
Prior studies report higher LLM test compilability, but typically use configurations that simplify the task through granularity restrictions, import handling, or repair. Siddiq et al. (2024a) generate one class per focal method and apply deterministic clean-up heuristics (stripping prose, fixing packages, bracket trimming), substantially inflating post-fix compilability; their raw SF110 compilability is very low with coverage under 2% without repairs. Yang et al. (2024) normalize imports and recursively prune failing test methods until classes compile, measuring coverage only on compilable residue; they attribute low validity to unresolved symbols, parameter mismatches, and abstract instantiation. Tang et al. (2024) achieve  70% runnable tests after light IDE fixes with a single prompt, yet EvoSuite dominates coverage. Yuan et al. (2024) document modest raw ChatGPT compilability (57.9% compilation errors) but large gains with iterative compile-feedback (ChatTester). We evaluate full classes without import injection or post-hoc repair, yielding lower compilability but exposing true failure mode distributions—including occasional EvoSuite failures on Defects4J. Our RQ2 confirms and refines prior taxonomies from Yang et al. (2024): beyond unresolved symbols and abstract instantiation, we distinguish cannot-find-symbol, missing-package, private-access, and redundant-constructors, quantifying how their prevalence shifts across CMD, Defects4J, and SF110. Uniquely, we analyze these errors under reasoning-oriented prompting, observing that such prompts reduce specific error classes (e.g., private-access violations) while leaving others (e.g., unresolved symbols) largely unaffected. Our findings extend broader code-generation hallucination accounts to unit-test generation, showing prompt design can selectively mitigate validity failures.
Summary of RQ2: LLMs generate a high proportion of syntactically correct tests ($\approx$90%), yet compilability remains low (<<10%) due to unresolved symbols, missing packages, and access violations. EvoSuite, while more reliable, also exhibits rare compilation failures ($\approx$1.9% in Defects4J), showing that errors are not unique to LLMs. Prompt engineering plays a critical role: reasoning-based prompts (CoT, GToT) improve structural correctness and reduce certain hallucinations, whereas FSL consistently underperforms. Compared to prior studies that operate at the method level with imports and repairs, our class-level, no-repair setup highlights the additional challenges of generating fully compilable suites directly from raw LLM outputs.
4.4 [RQ3]: Static Analysis of Human-Oriented Quality
[Experiment Goal]: This research question evaluates the quality of LLM-generated test suites from a static analysis perspective. Unlike EvoSuite, which primarily targets structural coverage, here the focus is on human-oriented dimensions such as coding standard compliance, maintainability (via cyclomatic and cognitive complexity), and potential defects. By analyzing these aspects, we assess whether prompt design can improve the readability and long-term usability of automatically generated tests.
[Experiment Design]: We evaluate test quality using three approaches. Checkstyle assesses adherence to Google Java Style Checkstyle and Sun Code Conventions, measuring formatting consistency and standard violations. PMD analyzes cyclomatic complexity (independent paths) and cognitive complexity (comprehension ease) (Dantas and Maia, 2021). SpotBugs detects potential defects categorized by severity: Scariest (critical issues), Scary (high-risk errors), Troubling (moderate concerns), and Of Concern (minor warnings). We compared prompting techniques (ZSL, FSL, CoT, ToT, GToT) across SF110, Defects4J, and CMD datasets.
@@
Coding Standard Compliance. LLM-generated test suites show varying coding standard adherence (Figure 6). FSL achieves highest compliance with Google Java Style and Sun Code Conventions, producing fewer violations overall.
(a) Heatmap of Google Code Style Violations
(b) Heatmap of Sun Code Style Violations
Figure 6: Heatmap of the Top-5 Google and Sun Code Style Violations Across Prompt Engineering
For Google Java Style (Figure 6(a)), indentation violations dominate across all prompting techniques, with GToT showing the lowest count (1757.5 vs. 2305.9 for ZSL). However, GToT increases import-related violations (140.6 CustomImportOrder and 93.1 AvoidStarImport violations) compared to FSL (71.3 and 23.4 respectively). For Sun Code Conventions (Figure 6(b)), MissingJavadocMethod violations are highest with CoT (566.5) and GToT (564.0), while GToT significantly reduces MagicNumber violations (134.8 vs. 323.4 for ZSL). These patterns suggest that reasoning-based prompts (CoT, ToT, GToT) improve code structure but sacrifice documentation and import organization. FSL strikes the best balance, minimizing both structural and documentation violations. These findings partially align with  Tang et al. (2024), who reported high ZSL violations, though our results show FSL and CoT significantly improve compliance while ToT and GToT produce tests with better logical structure but less style adherence.
Finding 13: FSL achieves the highest coding standard compliance, followed by CoT, which also reduces violations and improves adherence to style guidelines. In contrast, GToT enhances logical organization but introduces more violations, reinforcing the trade-off between structured prompting and strict adherence to coding conventions.
Code Complexity Analysis. Table 11 shows that all prompting strategies produce test suites with very low structural complexity: more than 99.5% of outputs fall into the lowest cyclomatic and cognitive complexity bands. GToT shows slightly higher proportions of simple cases (99.99% cyclomatic, 100% cognitive) and is the only strategy that avoids all moderate or high cases in cognitive complexity while nearly eliminating them in cyclomatic complexity (0.01% moderate, 0% high). Although these differences are marginal in absolute terms, they point to a stabilizing effect of GToT, reducing structural anomalies rather than changing the overall average complexity.
@@
These findings refine the observations of Tang et al. (2024), who showed that ZSL already produces simple tests, by highlighting that reasoning-based prompting can maintain this simplicity while reducing occasional outlier cases.
Finding 14: All prompting strategies yield predominantly low-complexity test suites. GToT achieves slightly higher proportions of simple outputs and is the only strategy with zero moderate or high cases in cognitive complexity, and nearly none in cyclomatic complexity. While differences across strategies are small, GToT appears to offer more structural consistency by reducing complexity outliers.
Static Analysis with SpotBugs.
Table 12 reports the distribution of static analysis warnings (SpotBugs) across 20,530 compilable test suites. While most defects are classified as minor “Of Concern” (55.6%), we observe notable differences across prompting strategies. FSL produces the most “Scariest” defects (e.g., 410 on SF110), while GToT yields the highest number of “Scary” (825) and “Troubling” (825) issues, particularly on SF110. These results highlight a concerning trade-off: although GToT improves structure and readability, it tends to introduce more subtle but structurally embedded defects. ZSL, while producing simpler tests, yields far fewer severe warnings. Unlike prior work (Tang et al., 2024), which focused on ZSL with limited compilation filtering, our evaluation reveals how reasoning-oriented prompting can unintentionally lead to more severe code issues, emphasizing the need for robust post-generation validation.
@@
ZSL = Zero-shot Learning, FSL = Few-shot Learning, CoT = Chain-of-Thought, ToT = Tree-of-Thought, GToT = Guided Tree-of-Thought. SpotBugs ranks issue severity as follows: Scariest (most critical), Scary, Troubling, and Of Concern (least critical).
Finding 15: Prompting strategies impact bug prevalence differently. FSL introduces the most critical (“Scariest”) issues, while GToT yields the highest number of structurally embedded “Scary” and “Troubling” bugs, particularly on SF110. ZSL produces the fewest severe warnings, albeit with simpler outputs. These results call for stricter validation when using reasoning-based prompting.
Relation to prior work.
Tang et al. (2024) analyze human-oriented quality through Checkstyle, PMD, and SpotBugs, showing that ChatGPT-generated tests under zero-shot prompting often violate style guidelines but remain structurally simple, with low cyclomatic and cognitive complexity. Yuan et al. (2024) complement this perspective with a user study, where developers judged LLM-generated tests comparable to manual ones in terms of usability and adoption effort. By contrast, neither Yang et al. (2024) nor Siddiq et al. (2024a) evaluate readability or maintainability, focusing instead on correctness and coverage. Our study extends this line of work by systematically comparing prompting strategies (ZSL, FSL, CoT, ToT, GToT) and showing how they shift coding standard compliance, complexity, and defect incidence. Unlike prior studies limited to zero-shot prompting, we demonstrate that reasoning-based prompts can improve logical organization at the cost of higher defect density, while FSL maximizes style compliance. This nuanced trade-off analysis situates prompt engineering as a key determinant of human-oriented quality in LLM-generated tests, beyond what was captured in earlier evaluations.
Summary of RQ3: LLM-generated tests generally exhibit low complexity and readable structures, consistent with prior reports. However, coding-standard compliance and defect incidence vary strongly with prompting strategies. FSL yields the cleanest style adherence, while reasoning-based prompts (CoT, ToT, GToT) enhance logical organization but introduce more subtle bugs flagged by SpotBugs. Compared to prior work limited to zero-shot prompting, our analysis shows that prompt engineering shifts the balance between readability, maintainability, and defect-proneness, underscoring its central role in shaping human-oriented quality.
@@
[Experiment Goal]: This experiment investigates whether LLM-generated test suites meet developer expectations of readability, focusing on clarity, structure, and ease of navigation. We use the readability model by Scalabrino et al. (2018), which has been validated against human judgments, to quantify readability beyond surface-level formatting. EvoSuite is included here as a baseline reference: although it does not explicitly optimize for readability, its widespread use in SBST makes it a valuable point of comparison to assess whether LLM prompting strategies deliver tangible improvements in human-oriented quality.
[Experiment Design]: We employ the readability model of Scalabrino et al. (2018), trained on Java code including JUnit tests and validated against human judgments. The model evaluates syntactic structure, identifier clarity, textual coherence, and overall structural readability, thus going beyond superficial formatting checks. All analyses are restricted to compilable test suites, ensuring that readability is measured on executable outputs. The evaluation is conducted in three complementary stages. First, a dataset-wise analysis: we aggregate readability scores by dataset (Defects4J, SF110, CMD) for GPT-3.5-turbo and Mistral 7B. EvoSuite is also included on Defects4J and SF110, where it was executable, but excluded from CMD due to Java version incompatibilities (Section 3.6). This view highlights how readability varies with dataset characteristics. Second, a prompt-wise analysis: we aggregate results by prompting strategy (ZSL, FSL, CoT, ToT, GToT) for GPT-3.5-turbo and Mistral 7B, and contrast them with EvoSuite on Defects4J and SF110. Although EvoSuite does not explicitly optimize for readability, its inclusion provides a widely used SBST baseline against which to assess the added value of prompting strategies. Finally, for CMD we restrict the analysis to LLMs only, since EvoSuite could not be executed due to Java version incompatibilities. All LLMs are evaluated across the five prompting strategies, but readability scores are reported only for configurations that yielded compilable test suites. For every configuration (dataset-wise, prompt-wise, CMD-only), readability distributions are summarized using descriptive statistics (minimum, quartiles, median, maximum, mean). Higher scores indicate more structured, navigable, and human-readable test suites.
@@
Dataset-Level Readability Analysis.
Table 13 shows LLM-generated test suites consistently achieve higher readability than EvoSuite across all datasets. Readability varies with codebase complexity: Defects4J exhibits highest readability scores (GPT-3.5-Turbo: 81.73%, Mistral 7B: 69.06% vs. EvoSuite: 57.33%), followed by SF110 (GPT-3.5-Turbo: 75.95%, Mistral 7B: 70.28% vs. EvoSuite: 50.67%). CMD presents lowest scores (GPT-3.5-Turbo: 60.93%, Mistral 7B: 57.13%), indicating modern, evolving codebases challenge LLMs more. While LLMs consistently outperform EvoSuite in readability, their effectiveness varies with dataset complexity.
Finding 16: LLM-generated tests outperform EvoSuite in readability. However, readability varies with dataset complexity, as shown in Table 13. CMD, which contains modern and evolving codebases, presents the greatest challenge, yielding the lowest readability scores across all models.
@@
  • 1
    Notice. Values are aggregated over all prompting strategies that produced compilable test suites.
LLM vs. EvoSuite: Readability Performance.
Table 14 shows LLM-generated tests consistently surpass EvoSuite in readability. For GPT-3.5-Turbo, GToT achieves the highest score (84.75% vs. EvoSuite’s 50.77%), followed by CoT (82.77%) and ToT (73.32%), while ZSL (77.93%) and FSL (73.49%) score lower. Similarly, for Mistral 7B, GToT leads (76.35%), followed by ToT (69.64%), with ZSL (63.79%) and FSL (60.40%) producing lowest scores. These results demonstrate reasoning-based prompting’s importance for readability.
Finding 17: LLM-generated test suites improve readability over EvoSuite by 21–40%, as shown in Table 14. This highlights the significant advantage of LLMs in producing more comprehensible test code.
@@
Prompting Strategies: Influence on Readability.
Table 15 shows reasoning-based prompting significantly enhances test suite clarity. GToT and CoT consistently achieve highest readability scores across datasets. For GPT-3.5-Turbo on CMD, CoT leads (67.99%), followed by GToT (61.86%), while ZSL (57.67%) and FSL (55.76%) score lower. Similarly, for GPT-4 on CMD, GToT achieves highest readability (67.71%). These results confirm minimal prompting provides insufficient guidance for well-structured tests, while reasoning-based approaches are essential for producing maintainable test suites.
@@
Finding 18: CoT and GToT consistently enhance readability, even on complex datasets like CMD, where unstructured prompting results in lower scores.
Model-Specific Readability Trends.
Table 15 reveals how different LLMs respond to reasoning-based prompting. GPT-3.5-Turbo performs best with CoT (67.99%), while GPT-4 excels with GToT (67.71%), confirming larger models benefit from structured reasoning. Mistral 7B shows higher variability, reaching 64.61% with GToT, indicating sensitivity to prompt design. Mixtral 8x7B performs exceptionally with ToT (79.93%), though its limited sample size affects stability.
Finding 19: Model architecture, parameter size, and prompting strategy significantly impact readability, with larger models consistently improving when guided by reasoning-based prompting.
Relation to prior work.
Readability from a developer perspective has been explicitly studied by Yuan et al. (2024), who conducted a user study showing that ChatGPT-generated tests were rated comparable to manual ones in terms of readability and adoption effort. Other evaluations, such as Tang et al. (2024), focus on understandability proxies (e.g., style violations, complexity) rather than direct readability. Our study extends this line of work by systematically quantifying readability with an established model validated against human judgments, and by comparing LLMs with EvoSuite. We further show how prompting strategies (CoT, ToT, GToT) significantly enhance readability beyond zero-shot and few-shot settings, demonstrating that structured reasoning prompts align LLM outputs more closely with developer expectations.
Summary of RQ4: LLM-generated tests consistently achieve higher readability than EvoSuite, with gains of 21–40% across datasets. Readability varies with dataset complexity, being highest on Defects4J and lowest on CMD. Reasoning-based prompting (CoT, GToT) markedly improves clarity and navigability, while zero-shot and few-shot remain weaker. Compared to prior work, which assessed readability through targeted developer studies, our study provides a systematic, model-based evaluation showing that prompting strategies play a central role in aligning LLM outputs with developer expectations.
@@
[…]
@@
EvoSuite vs. LLMs: Overall Coverage Performance.
Tables 16, 17, and 18 reveal a significant performance gap between LLMs and EvoSuite. EvoSuite consistently outperforms LLMs across all metrics, achieving median coverage of 94.34% (line), 95.0% (instruction), and 100% (method) through search-based optimization. LLM-generated tests show higher variability, with reasoning-based prompting significantly improving coverage over ZSL and FSL. GPT-3.5-Turbo with CoT achieves the highest LLM median coverage (75.47% line, 76.0% instruction, 100% method), though all LLM configurations fall short of EvoSuite’s systematic exploration capabilities.
@@
Finding 20: EvoSuite outperforms all LLMs in coverage metrics. However, reasoning-based techniques (CoT, GToT, ToT) significantly enhance LLM test effectiveness, reducing the performance gap.
Dataset-Level Coverage Trends.
Tables 1617, and 18 show coverage varies significantly across datasets. Defects4J achieves highest coverage (GPT-3.5-Turbo: 82.74% line, 85.0% instruction, 100% method), suggesting its structured design facilitates test generation with clearer functional boundaries. In contrast, SF110 reports much lower coverage (3.57% line, 2.0% instruction, 16.67% method) due to higher complexity and diverse architectures. CMD reports highest median coverage (100% for some prompts), though this likely reflects its smaller size and modular structure rather than superior LLM capabilities. The contrast between datasets highlights structural influences: Defects4J benefits from potential training exposure (Sallou et al. (2023)), clearer functional boundaries, and modularized code, while SF110 presents challenges through higher complexity and lower public accessibility.
Finding 21: LLMs achieve higher coverage on well-documented datasets like Defects4J but struggle with complex projects like SF110, where greater structural variability makes test generation more challenging. CMD’s high coverage is likely due to its small, modular dataset rather than an inherent LLM advantage.
@@
Finding 22: Dataset characteristics influence LLM performance—Defects4J’s high coverage likely stems from pretraining exposure while SF110’s limited representation restricts generalization, highlighting the need for careful dataset selection to mitigate data leakage in test generation studies.
Impact of Reasoning-Based Prompting on Coverage.
Tables 161718 show reasoning-based techniques (CoT, GToT, ToT) consistently improve coverage compared to ZSL and FSL.
For line coverage, GPT-3.5-Turbo with CoT achieves highest median (75.47%), while ZSL and FSL produce significantly lower coverage (27.68% and 16.14%). Mistral 7B performs best with GToT (30.0%), though still below GPT-3.5-Turbo.
For instruction coverage, EvoSuite maintains advantage (95.0%). Among LLMs, GPT-3.5-Turbo with CoT (76.0%) and GToT (45.0%) lead, while Mistral 7B’s best result with GToT (26.0%) indicates smaller models’ limitations.
For method coverage, GPT-3.5-Turbo with CoT reaches 100% in some cases, while Mistral 7B improves with GToT (89.59%) but still lags. These results demonstrate reasoning-based prompting enhances LLM performance significantly but doesn’t fully bridge the gap with EvoSuite.